# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

### v1

@@ 2.1 Least squares
Thus, surprisingly, we demonstrate that the least squares estimate for the unknown true bias-correction term $h_{0}$ can be defined by an objective function that does not explicitly include $h_{0}$ (or the unknown $r_{0}$) itself. As discussed in the following subsection, this objective function can be easily approximated using observations.
2.2 Empirical risk minimization
We then estimate $r_{0}$ by solving the following empirical risk minimization:
$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
where the empirical risk is given as
$\widehat{{\mathcal{L}}}_{n}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r(1,X_{i})-2r(0,X_{i})+\mathbbm{1}[D_{i}=1]r(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r(0,X_{i})^{2}\Big),$
and $J\colon{\mathcal{R}}\to{\mathbb{R}}^{+}$ is a regularization term with a penalty coefficient $\lambda>0$.
This type of estimation method is referred to as least-squares importance fitting (LSIF, Kanamori et al., 2009) in the literature on density-ratio estimation (DRE).
@@ 3 Estimation error analysis
This section provides an estimation error analysis for $r_{0}$ estimated by the direct bias-correction term estimation method. For simplicity, throughout the analysis, we use the following model for $e_{0}$: $e(x)=\frac{1}{1+\exp(-f(x))}$, where $f\colon{\mathcal{X}}\to{\mathbb{R}}$ is a function belonging to a set $\mathcal{F}$}. We use various models for $\mathcal{F}$}, including linear models, RKHS, and neural networks. Then, the model of $r_{0}$ is given as $r(1,x)=1+\exp(-f(x)),\quad r(0,x)=\exp(f(x))$. Let $f_{0}$ be the true function for $e_{0}$; that is, $e_{0}(X)=\frac{1}{1+\exp(-f_{0}(x))}$.
3.1 Linear models
First, we consider the case where $f_{0}$ belongs to a linear model. There exists $\theta_{0}\in\Theta^{K}$ such that $e_{0}(x)=\frac{1}{1+\exp\big(-x^{\top}\theta_{0}\big)}$, where $\Theta$ is a compact parameter space. Under this assumption, we have $r_{0}(1,X)=1+\exp\big(-x^{\top}\theta_{0}\big)$ and $r_{0}(0,X)=1+\exp\big(x^{\top}\theta_{0}\big)$. Let us use $r_{\theta}(1,X)=1+\exp\big(-x^{\top}\theta\big)$ and $r_{\theta}(0,X)=1+\exp\big(x^{\top}\theta\big)$ as a model for $r_{0}$. Then, we define the estimator as $\widehat{r}^{\mathrm{Lin}}_{n}=r_{\widehat{\theta}_{n}}$, where $\widehat{\theta}_{n}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$. Here, we set $\lambda=0$.
We show the asymptotic normality of this estimator.
@@ 3.1 Linear models
We assume the following: (i) $\theta_{0}$ is in the interior of $\Theta$; (ii) $\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$ is twice continuously differentiable on some neighborhood $\mathcal{M}$} of $\theta_{0}$ (with probability one); (iii) $\sqrt{n}\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})\xrightarrow{d}{\mathcal{N}}(0,\Sigma_{0})$; (iv) for any sequence $\tilde{\theta}_{n}\xrightarrow{p}\theta_{0}$, it holds that $\frac{\partial^{2}}{\partial\theta\partial\theta^{T}}\widehat{{\mathcal{L}}}_{n}(r_{\theta})-B_{0}\xrightarrow{p}0$, for some non-stochastic $K\times K$ matrix $B_{0}$ that is nonsingular.
Then, from the asymptotic theory of extremum estimators, we obtain the following result:
Lemma 3.2.
Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{\theta}_{n}-\theta_{0}\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,B_{0}^{-1}\Sigma_{0}B_{0}^{-1})$.
From the Taylor expansion of $r_{\theta}(x)$ around $\theta=\theta_{0}$, Lemma 3.2 immediately yields the following result:
Theorem $3.$
Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{r}^{\mathrm{Lin}}_{n}(x)-r_{0}(x)\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,\kappa(x))$, where $\kappa(x)\coloneqq\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)^{\top}B_{0}^{-1}\Sigma_{0}B_{0}^{-1}\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)$.
3.2 RKHS
Next, we investigate the case with RKHS regression. Let ${\mathcal{R}}^{\mathrm{RKHS}}$ be a class of RKHS functions, and define the estimator as $\widehat{r}^{\mathrm{RKHS}}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}^{\mathrm{RKHS}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda\|r\|^{2}_{{\mathcal{H}}}$, where $\|\cdot\|^{2}_{{\mathcal{H}}}$ is the RKHS norm. We analyze the estimation error by employing the results in Kanamori et al. (2012), which study RKHS-based LSIF in DRE.
We define the following localized class of RKHS functions as a technical device: ${\mathcal{R}}^{\text{RKHS}}_{M}\coloneqq\big\{r\in{\mathcal{R}}^{\text{RKHS}}\colon I(r)\leq M\big\}$. We then make the following assumption using this localized class.
Assumption 3.4.
There exist constants $0<\gamma<2$, $0\leq\beta\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{R}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
@@ 3.2 RKHS
Suppose that Assumptions 1.1 and 3.4 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, then we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|\big(1-e_{0}(X)\big)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=O_{P_{0}}\left(\lambda^{1/2}\right)$.

### v2

@@ 2.1 Population Bregman divergence minimization
Thus, surprisingly, we demonstrate that the least squares estimate for the unknown true bias-correction term $h_{0}$ can be defined by an objective function that does not explicitly include $h_{0}$ itself. As discussed in the following subsection, this objective function can be easily approximated using observations.
2.2 Empirical Bregman divergence minimization
Then, we estimate the bias-correction term $h_{0}$ by minimizing an empirical Bregman divergence as
$\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g}\big(h\big)+\lambda J(h),$
where $J(h)$ is some regularization function and
$\widehat{\mathrm{BR}}_{g}(h)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(h(D_{i},X_{i}))+\partial g(h(D_{i},X_{i}))h(D_{i},X_{i})-\partial g(h(1,X_{i})+\partial g(h(0,X_{i}))\Big).$
2.3 Losses for the bias-correction term estimation
By changing $g$, we can obtain various loss functions for estimating the bias-correction term, as shown in the subsequent subsections. In particular, if we use the squared loss in the Bregman divergence, we obtain Riesz regression in Chernozhukov et al. (2021), which is originally called Least-Squares Importance Fitting (LSIF) in the DRE literature Kanamori et al. (2009). Note that kernel mean matching by Gretton et al. (2009) is also the same as, or a variant of, LSIF. If we use the KL divergence, we obtain the tailored loss in Zhao (2019), which is originally called KLIEP in the DRE literature Sugiyama et al. (2008). Furthermore, as we discuss in Section 3, if we use linear models for $h_{0}$ and train them with the squared loss, the covariate balancing property is automatically obtained, as shown in Bruns-Smith et al. (2025). If we model $h_{0}$ using the propensity score $e_{0}$ approximated via logistic models and train it with the tailored loss, the covariate balancing property is automatically obtained, as shown in Zhao (2019). We demonstrate the correspondence of the existing methods in Table 1. Also see Figure 1 in Appendix for the relationship among bias-correction term estimation via Bregman divergence minimization, density ratio estimation, and covariate balancing, summarized in Kato (2025a) and Kato (2025c).
Table 1: Correspondence among DRE methods and bias-correction term estimation methods (BCE).
DRE method BCE method $g(t)$
LSIF (Kanamori et al., 2009) Riesz regression (Chernozhukov et al., 2021) $(t-1)^{2}/2$
Kernel Mean Matching (Gretton et al., 2009) Stable balancing weights (Zubizarreta, 2015)
UKL (Nguyen et al., 2010) Tailored loss (Zhao, 2019) $t\log(t)-t$
KLIEP (Sugiyama et al., 2008) Entropy balancing weights (Hainmueller, 2012)
Binary KL divergence $t\log(t)-(1+t)\log(1+t)$
PULogLoss (Kato et al., 2019) $C\log\left(1-t\right)$
$Ct\left(\log\left(t\right)-\log\left(1-t\right)\right)$ for $0<t<1$
2.4 Squared loss
Our least squares method for direct bias-correction term estimation can be obtained by using a squared loss $g^{\text{SL}}(h)=(h-1)^{2}$. By substituting this function into the Bregman divergence, we formulate the estimation problem as $h^{*}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g^{\text{SL}}}\big(h\big)$, where
$\mathrm{BR}_{g^{\text{SL}}}\big(h\big)={\mathbb{E}}\Big[-2\big(h(1,X)-h(0,X)\big)+h(D,X)^{2}\Big].$
Then, we estimate the bias-correction term as $\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g^{\mathrm{SL}}}\big(h\big)+\lambda J(h)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{SL}}}(h)=\frac{1}{n}\sum^{n}_{i=1}\big(-2\big(h(1,X_{i})-h(0,X_{i})\big)+h(D_{i},X_{i})^{2}\big)$. This objective function is the same as the one used in Chernozhukov et al. (2021). This type of estimation method is referred to as LSIF in density-ratio estimation (Kanamori et al., 2009).
2.5 KL divergence loss
Consider $g^{\mathrm{KL}}(h)=|h|\log|h|-|h|$, which is a convex function. By substituting this function into the Bregman divergence, we formulate the estimation problem as $h^{*}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g^{\mathrm{KL}}}\big(h\big)$, where
$\mathrm{BR}_{g^{\mathrm{KL}}}\big(h\big)\coloneqq{\mathbb{E}}\Big[|h(D_{i},X_{i})|-\log(|h(1,X)|)-\log(|h(0,X)|)\Big].$
Then, we estimate the bias-correction term as $\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g^{\mathrm{KL}}}\big(h\big)+\lambda J(h)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{KL}}}(h)=\frac{1}{n}\sum^{n}_{i=1}\big(|h(D_{i},X_{i})|-\log\big(|h(1,X_{i})|\big)-\log\big(|h(0,X_{i})|\big)\big)$. This estimation method corresponds to unnormalized Kullback–Leibler (UKL) minimization in DRE (Nguyen et al., 2010), which generalizes the KL importance estimation procedure (KLIEP). Also see Appendix B.
2.6 Tailored loss (a variant of the KL divergence loss)
Next, as a variant of the KL divergence loss, we propose the tailored loss. Let us redefine a model $\mathcal{H}$} as a set of functions $h(1,\cdot)\colon{\mathcal{X}}\to(1,\infty)$ and $h(0,\cdot)\colon{\mathcal{X}}\to(-1,-\infty)$; that is, we restrict the space of $h$. This restriction is justified from the form of $h_{0}$ and the common support assumption. Let us consider $g^{\text{TL}}(h)=(|h|-1)\log\left(|h|-1\right)-|h|$. By substituting this function, we obtain
$\mathrm{BR}_{g^{\text{TL}}}\big(h\big)\coloneqq{\mathbb{E}}\Big[\log\left(|h(D,X)|-1\right)+|h(D,X)|-\log\left(|h(1,X)|-1\right)-\log\left(|h(0,X)|-1\right)\Big].$
Note that it holds that $\mathrm{BR}_{g^{\text{TL}}}\big(h\big)\coloneqq{\mathbb{E}}\Big[-\mathbbm{1}[D=0]\log\left(|h(1,X)|-1\right)-\mathbbm{1}[D=1]\log\left(|h(0,X)|-1\right)+\mathbbm{1}[D=1]h(1,X)-\mathbbm{1}[D=0]h(0,X)\Big]$. Then, we estimate the bias-correction term as $\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g^{\text{TL}}}\big(h\big)$, where the empirical Bregman divergence becomes $\widehat{\mathrm{BR}}_{g^{\text{TL}}}(h)=\frac{1}{n}\sum^{n}_{i=1}\big(\mathbbm{1}[D_{i}=0]\log\big(|h(1,X_{i})|-1\big)+\mathbbm{1}[D_{i}=1]\log\big(|h(0,X_{i})|-1\big)+\mathbbm{1}[D_{i}=1]|h(1,X_{i})|-\mathbbm{1}[D_{i}=0]|h(0,X_{i})|\big)$.
3 Automatic covariate balancing
Under specific choices of Riesz regression models and Bregman divergence, we can automatically enforce the covariate balancing property. The key tool is the duality relationship between the Bregman divergence minimization problem and the covariate balancing optimization problem. This result is shown in Kato (2025a), and we introduce the result for reference.
3.1 Linear models and squared loss
Consider a linear model
$h_{\beta}(D,X)=\Phi(D,X)^{\top}\beta,$
where $\Phi\colon\{1,0\}\times{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. For this model, using the squared loss (Riesz regression) automatically achieves covariate balancing, as discussed in Bruns-Smith et al. (2025).
Specifically, under linear models, by duality, this MSE minimization problem is equivalent to solving
$\min_{w\in{\mathbb{R}}^{n}}\|w\|^{2}_{2}\quad\text{s.t.}\ \ \sum^{n}_{i=1}w_{i}\Phi(D_{i},X_{i})-\left(\sum^{n}_{i=1}\Big(\Phi(1,X_{i})-\Phi(0,X_{i})\Big)\right)=\bm{0}_{p},$
where $\bm{0}_{p}$ is the $p$-dimensional zero vector. This optimization problem matches that used to obtain stable weights (Zubizarreta, 2015).
It enforces the covariate balancing condition $\sum^{n}_{i=1}\widehat{w}_{i}\Phi(D_{i},X_{i})-\left(\sum^{n}_{i=1}\Big(\Phi(1,X_{i})-\Phi(0,X_{i})\Big)\right)=\bm{0}_{p}$, where $\widehat{w}_{i}=\Phi(D_{i},X_{i})^{\top}\widehat{\beta}$.
Another advantage of using linear models is that we can write the entire ATE estimation with a single linear model, as shown by Bruns-Smith et al. (2025).
3.2 Logistic models and tailored loss
We can model the Riesz representer by modeling the propensity score as
$h_{\beta}(D,X)=\mathbbm{1}[D=1]r_{\beta}(1,X)-\mathbbm{1}[D=0]r_{\beta}(0,X),$
where $r_{\beta}(1,X)=\frac{1}{e_{\beta}(X)}$, $r_{\beta}(0,X)=\frac{1}{1-e_{\beta}(X)}$, $e_{\beta}(X)\coloneqq\frac{1}{1+\exp\big(-\beta^{\top}\Phi(X)\big)}$, and $\Phi\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. Note that we do not include $D$, unlike the basis function used in linear models. For this model, if we use the KL-divergence–flavored convex function defined in Section 2.6, which corresponds to the tailored loss in Zhao (2019), we automatically achieve covariate balancing.
Define $\widehat{\beta}\coloneqq\operatorname*{arg\,min}_{\beta}\frac{1}{n}\sum^{n}_{i=1}\sum_{d\in\{1,0\}}\Big(\mathbbm{1}[D_{i}=d]\left(-\log\left(\frac{1}{r_{\beta}(d,X_{i})-1}\right)+r_{\beta}(d,X_{i})\right)\Big)$, and denote $r_{\widehat{\beta}}$ by $\widehat{r}$. Under logistic models, by duality, the KL divergence-flavored loss is equivalent to solving
$\min_{w\in(1,\infty)^{n}}\sum^{n}_{i=1}(w_{i}-1)\log(w_{i}-1)\quad\text{s.t.}\ \ \left(\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]w_{i}\Phi(X_{i})-\mathbbm{1}[D_{i}=0]w_{i}\Phi(X_{i})\Big)\right)=\bm{0}_{p}.$
This optimization problem matches that used in entropy balancing (Hainmueller, 2012). Note that this objective function is derived from $\widehat{\mathrm{BR}}_{g^{\text{TL}}}(h)$ when we use the logistic model specified in this section.
As a result, we obtain $\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\widehat{w}_{i}\Phi(X_{i})-\mathbbm{1}[D_{i}=0]\widehat{w}_{i}\Phi(X_{i})\Big)=\bm{0}_{p}$, where $\widehat{w}_{i}=\widehat{r}(X_{i}).$
This model has the advantage that we can use a basis function $\Phi(X)$ independent of $D$. Moreover, it naturally achieves covariate balance in the sense that the covariate distributions match between the treated and control groups. Additionally, it allows us to automatically impose nonnegativity on $h(1,X)$ and h(0,X)-h(0,X), which may be violated in linear models. Note that $h_{0}(1,X)=\frac{1}{e(X)}$ and $h_{0}(1,X)=\frac{1}{1-e(X)}$.
3.3 Comparison
We first discuss the advantages of using logistic models over linear models. One benefit of using logistic models is that we can simplify the basis function by making it independent of $D$. Furthermore, we can express covariate balancing in a clearer form as $\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\widehat{w}_{i}\Phi(X_{i})-\mathbbm{1}[D_{i}=0]\widehat{w}_{i}\Phi(X_{i})\Big)=\bm{0}_{p}$, while under linear models, $\sum^{n}_{i=1}\widehat{w}_{i}\Phi(D_{i},X_{i})-\left(\sum^{n}_{i=1}\Big(\Phi(1,X_{i})-\Phi(0,X_{i})\Big)\right)=\bm{0}_{p}$ is attained, but it is somewhat harder to interpret. Moreover, using logistic models incorporates more information about the form of the bias-correction term, which includes the inverse propensity function. Logistic models also naturally impose restrictions such that $h(1,X)\in(1,\infty)$ and $h(0,X)\in(-\infty,-1)$ under the common support assumption.
In contrast, if we use linear models, we can express the entire ATE estimator with a single linear model, as shown in Bruns-Smith et al. (2025). Furthermore, we can obtain the estimator of the bias-correction term as a closed-form solution. In addition, as discussed in Kato (2025b), a subsequent work of this study. nearest neighbor matching is also an instance of linear models trained via Riesz regression (squared loss). We introduce the result in Appendix H for reference.
Ultimately, there is no clear dominance between the use of linear and logistic models. Moreover, we can also use more complex models, such as random forests and neural networks. The choice of model should be made based on the data and application, and once the model is selected, we can determine appropriate specifications that ensure covariate balancing automatically.
@@ 4 Estimation error analysis
This section provides an estimation error analysis for $h_{0}$ estimated by the direct bias-correction term estimation method. We can use various models for $\mathcal{H}$}, including RKHS and neural networks.
4.1 Model
We define a model of the bias-correction term $h_{0}$ by $h(D,X)=\zeta^{-1}\circ f(D,X)=\zeta^{-1}(f(D,X))$, where $\zeta$ is a continuously differentiable and globally Lipschitz link function, and $f$ is some basic model. For example, if we use linear model for the bias-correction term $h_{0}$, we can write $h(D,X)=\Phi(D,X)^{\top}\beta$, where $\zeta$ is the identity function, $f(D,X)=\Phi(D,X)^{\top}\beta$, $\Phi$ is some basis function and $\beta$ is the corresponding parameter. If we use logistic model for the bias-correction term $h_{0}$, we can use logistic link for $\zeta$, and $f(D,X)=\Phi(X)^{\top}\beta$.
4.2 RKHS
First, we investigate the case with RKHS regression. Let ${\mathcal{F}}^{\mathrm{RKHS}}$ be a class of RKHS functions, and define $\widehat{f}^{\mathrm{RKHS}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{RKHS}}}\widehat{{\mathcal{L}}}_{n}(\zeta^{-1}\circ f)+\lambda\|f\|^{2}_{{\mathcal{F}}}$, where $\|\cdot\|^{2}_{{\mathcal{F}}}$ is the RKHS norm. Then, we define an estimator as $h^{\mathrm{RKHS}}\coloneqq\zeta^{-1}\circ\widehat{f}^{\mathrm{RKHS}}_{n}$ We analyze the estimation error by employing the results in Kanamori et al. (2012), which study RKHS-based LSIF in DRE. We define the following localized class of RKHS functions as a technical device: ${\mathcal{F}}^{\text{RKHS}}_{M}\coloneqq\big\{f\in{\mathcal{F}}^{\text{RKHS}}\colon I(f)\leq M\big\}$ for some norm $I(f)$ of $f$. We also define ${\mathcal{H}}^{\text{RKHS}}\coloneqq\left\{\zeta^{-1}\circ f\colon f\in{\mathcal{F}}^{\mathrm{RKHS}}\right\}$. We then make the following assumption using this localized class.
@@ 4.2 RKHS
There exist constants $0<\gamma<2$, $0\leq\beta\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{F}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
@@
Suppose that $g$ is $\mu$-strongly convex and there exist constant $C>0$ such that $|g^{\prime\prime}(t)|\leq C\quad\forall t\in{\mathbb{R}}$. Assume also that $\zeta^{-1}(0)$ is finite. Suppose that Assumptions 1.1 and 4.1 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $h_{0}\in{\mathcal{H}}^{\mathrm{RKHS}}$, then we have $\left\|\widehat{h}^{\mathrm{RKHS}}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{P_{0}}\left(\lambda^{1/2}\right)$.